OpenStudy (anonymous):

convert to logarithmic function 4^(-2)=1/16

OpenStudy (smurfy14):

log[4]1/16=-2

OpenStudy (smurfy14):

\[\log_{4}1/16= \]

OpenStudy (smurfy14):

\[\log_4 1/16=-2 \]

OpenStudy (anonymous):

thx